The minute hand of a wall clock is 14 cm long. In the time interval from 6 PM to 6 : 30 PM, the tip of the
minute hand has

Grade:12th Pass

1 Answers

Sudheesh Singanamalla
13 years ago

Question is incomplete .

the minute hand has moved through 6 * 5/2 = 30/2 = 15 degrees !

the minute hand covered a distance of 15/180* 3.14 * 14 = 3.6633 centimetres.

 

Please approve !